# The impact of local noise recorded at the ET candidate sites on the signal to noise ratio of CBC gravitational wave signals for the ET triangle configuration

### v5

@@ IV.1 Seismic
Figure 4: Seismic ASD displacement for the two candidate sites, TERZ in red and P2 in blue, are reported as a percentile (50th in solid line while 10th and 90th the lower and upper limit of the colored band). Peterson limits (NLNM in black dashed line and NHNM in black dotted line) and the self-noise of both sensors (dash-dot line) are also shown.
@@
Figure 4 shows the comparison of the percentiles for P2 (in blue) and TERZ (in red) with respect to the Peterson New Low and High Noise Model (NLNM in black dashed line and NHNM in black dotted line respectively) [86]. Above 6 Hz, the P2 data lie on the self-noise of the sensor, making apparent the quietness of the site. We will use these data in Section V as input parameters for the determination of the ET sensitivity curve.
IV.2 Astrophysical Sources
@@
The parameters for IMBH and BNS mergers are taken from the catalog [87] used to produce the results of [8], which considers a set of sources spanning one year between the mock dates 01-01-2030 and 12-31-2030. For what concerns BBH, the catalog results from the mixture of BBH from isolated binary evolution, dynamical formation in young, globular and nuclear star clusters. Masses, spins, redshifts and luminosity distances have been obtained with the open-source code FASTCLUSTER [88]. For this work, we select the events that satisfy the following conditions:
  • BBH with at least one source frame component mass above 100 $M_{\bigodot}$;
  • BBH with both source frame component mass above 100 $M_{\bigodot}$;
  • BBH with a total mass above 100 $M_{\bigodot}$.
for a total of $\simeq 10^{3}$ events. Figure 5(a) shows the distribution of the time to the merger at $10\rm\,Hz$ for these events. We note that once left the frequency band considered in this work, most events have less than $0.5\rm\,s$ before merging. On the other hand Figure 5(b) shows the time spent in the $[2,10]\rm\,Hz$ range and points out how most of the time IMBH lie in this bandwidth before the merger.
((a)) Distribution of the time to merger at $10\rm\,Hz$ for the BBH events considered in this work.
@@
((b)) Distribution of the time spent in the $[2,10]\rm\,Hz$ range for the BBH events considered in this work.
Figure 5: Times distributions for the IMBH of this work.
((a)) Distribution of the time to merger at $10\rm\,Hz$ for the BNS events considered in this work.
((b)) Distribution of the time spent in the $[2,10]\rm\,Hz$ range for the BNS events considered in this work.
Figure 6: Times distributions for the BNS of this work.
@@
For BNS mergers, the source frame masses of the two objects are sampled uniformly in the interval $[1.1;2.5]\rm M_{\bigodot}$ so that $m_{1}>m_{2}$. Redshifts and luminosity distances are taken from [89]. Since [87] contains $\mathcal{O}(7\times 10^{5})$ BNS events, to limit the computational cost and time of the analysis we randmly sample, using a logarithmic distribution in redshift to privilege closer - therefore stronger - signals, $10^{3}$ events at $z<1.5$. As mentioned, the BNS signal should reach $SNR\geq 12$ to issue an early warning between 20 hours and 1 hour before the merger. Figure 6(a) shows the distribution of the time to the merger at $10\rm\,Hz$ for the selected BNS events. We note that, once left the [2,10] Hz frequency band, most events have less than $20\rm\,{min}$ before merger. Figure 6(b) shows the time spent in the $[2,10]\rm\,Hz$ range. The relevance of this bandwidth is apparent, since several hours are spent here.
@@
V Modified ET curves
The ET sensitivity, modified according to what described in Section III, takes into account the local seismic component of the NN of the two candidate sites. In equation 5, we assume $\rho_{0}=2.7\cdot 10^{3}$ kg/m3 and $L=10$ km. Figure 7 reports the sensitivity curve using the NN from P2 and TERZ compared with the ET design noise budget as reported in [8] for the frequency range between 1 Hz and 100 Hz.
@@ V Modified ET curves
To quantify the modified ET sensitivity with respect to the designed one, we evaluated the NTR index defined in Sec. III. While Fig. 7 shows the level of the NN inside the ET sensitivity averaged over two years, the NTR, computed every $120\rm\,s$, gives information on how many times the modified sensitivity ($S_{n}(f)$) is above ($>1$) ore below ($<1$) the designed one ($S_{n,\rm{ET}}(f)$), i.e., information related to the glitchiness of the site. Focusing our attention on the frequency range between $f_{1}=2$ Hz and $f_{2}=10$ Hz, a value of NTR larger than one implies that, in the selected dataset, the contribution of the NN is limiting the ET sensitivity. The higher the NTR, the higher the probability of losing the GW signals. One has to remember that if the bandwidth of the GW signal is larger than $f_{2}$ it is possible to recover the signal since the NN is not significant at higher frequencies.
@@
Percentile P2 TERZ P2 TERZ P2 TERZ P2 TERZ P2 TERZ
@@
VI Impact on Signal to Noise Ratio
VI.1 Intermediate Mass Black Holes
Figures 10 and 11 show the SNR distributions for the IMBH case for the Sardinia and EMR candidate sites, respectively. The benchmark SNR distribution inferred using the design ET sensitivity curve is shown in green. The results are also summarized in Table 2. In both cases, we analyze the same 970 events. For the analysis and visualization, we also exclude all those events which have $\rm SNR<12$ in the best noise condition already. These events never reach the SNR = 12 threshold at 10 Hz also when the noise worsens and therefore are not relevant for the scope of this work. On the contrary, this cut guarantees that events eventually lost in the design case, but recovered in the best noise conditions, are retained. This can also highlight any possible improvement of the performances with respect to the expectations.
In Sardinia, on average, the 10th and 50th percentile cases show comparable results (Figures 10(a) and 10(b)). The SNR shows a gain of $\simeq 6\%$ and $\simeq 5\%$, respectively (Figures 10(d) and 10(e)). The non overlapped green histogram bin, on the left of the SNR = 12 threshold, highlights a marginal improvement ($\simeq 2\%$) in the fraction of recovered events for the 10th and 50th percentiles with respect to the design case. On the other hand, in the 90th percentile case the performance is on par with the design case: there is a marginal SNR loss ($\simeq 2.5\%$ Figure 10(f)) and the fraction of lost events is consistent with the forecast of the design performance (Figure 10(c)).
For the EMR site, the situation is different. The 10th percentile case is on par with the design performance, with the SNR showing a gain of $\simeq 2.5\%$ (Figures 11(a) and 11(d)). In the 50th percentile, the SNR distribution starts do deviate from the design case, with $5\%$ of events below threshold and a SNR loss $\simeq 12\%$ (Figures 11(b) and 11(e)). The 90th percentile shows, as expected, the worst performance with an average SNR loss $\simeq 37\%$ and $24\%$ of events with $SNR\leq 12$ in the considered bandwidth (Figures 11(c) and 11(f)). Table 2 also shows the mean of the ratio between the SNR of the events in Sardinia and the SNR of the events at EMR.
@@ VI.1 Intermediate Mass Black Holes
Case Sardinia EMR Sardinia EMR
10th percentile $+6\%$ $+2.5\%$ $-2\%$ $-1\%$ $0.96$
50th percentile $+5\%$ $-12\%$ $-2\%$ $4\%$ $0.83$
90th percentile $-2.5\%$ $-37\%$ $2\%$ $24\%$ $0.69$
@@
Table 2: Summary of the SNR performance for the Sardinia and EMR candidate sites in the case of IMBH signals. The negative figures under the events with SNR $<12$ highlight an improvement in the fraction of recovered events with respect to the design case.
((a)) Sardinia Nominal SNR distribution (green) and SNR distribution using the 10th percentile ET sensitivity (blue).
((b)) Sardinia nominal SNR distribution (green) and SNR distribution using the 50th percentile ET sensitivity (blue).
((c)) Sardinia nominal SNR distribution (green) and SNR distribution using the 90th percentile ET sensitivity (blue).
((d)) Sardinia SNR/SNRdesign using the 10th percentile ET sensitivity.
((e)) Sardinia SNR/SNRdesign using the 50th percentile ET sensitivity.
((f)) Sardinia SNR/SNRdesign using the 90th percentile ET sensitivity.
Figure 10: IMBH SNR distributions for the Sardinia candidate site. In the top row, the SNR = 12 threshold is marked with a black dashed line. In the bottom row, the black dashed line represents the equality between the SNR measured in the design case with the cases of the modified ET curves. The non overlapped green SNR bin under the SNR threshold highlights the fact that there are some events lost in the design case but recovered in the best possible recorded noise conditions.
((a)) EMR Nominal SNR distribution (green) and SNR distribution using the 10th percentile ET sensitivity (blue).
((b)) EMR Nominal SNR distribution (green) and SNR distribution using the 50th percentile ET sensitivity (blue).
((c)) EMR Nominal SNR distribution (green) and SNR distribution using the 90th percentile ET sensitivity (blue).
((d)) SNR/SNRdesign at EMR using the 10th percentile ET sensitivity.
((e)) SNR/SNRdesign at EMR using the 50th percentile ET sensitivity.
((f)) SNR/SNRdesign at EMR using the 90th percentile ET sensitivity.
Figure 11: IMBH SNR distributions for the EMR candidate site. In the top row, the SNR = 12 threshold is marked with a black dashed line. In the bottom row, the black dashed line represents the equality between the SNR measured in the design case with the cases of the modified ET curves. The non overlapped green SNR bin under the SNR threshold highlights the fact that there are some events lost in the design case but recovered in the best possible recorded noise conditions.
VI.2 Binary Neutron Stars
Figures 12 and 13 and Table 3 show the results obtained in the case of BNS signals. As for the previous results, we exclude events with SNR $<12$ in the best noise case. In general, the results are consistent with the IMBH case. In the 10th percentile case, we appreciate a $2\%$ improvement in the fraction of recovered events with respect to the design case. For the 10th, 50th and 90th percentile cases, we find $\simeq 0.05\%$ and $\simeq 3\%$ of events below threshold in the $[2,10]\rm\,Hz$ range, respectively. These figures are on par, if not slightly better, with respect to the design case (Table 3). In all cases, the ratios between the reconstructed SNR and the design SNR show only marginal differences (Figures 12(d), 12(e) and 12(f)).
On the other hand, the EMR site shows less promising performances. The 10th percentile case (Figures 13(a) and 13(d)) equals the figure of the benchmark case, with a SNR gain $\simeq 2\%$ and $1\%$ more events recovered than the design case. The 50th and 90th percentile show the worst performance and loose $\simeq 6.5\%$ and $\simeq 22\%$ of events, respectively (Figures 13(b) and 13(c)). For what concerns the SNR ratios (Figures 13(e) and 13(f)), the figures are consistent with the IMBH case and both sites show a loss which is in line with what has been already discussed in Section VI.1. Table 3 also shows the mean of the ratio between the SNR of the events in Sardinia and the SNR of the events at EMR.
@@ VI.2 Binary Neutron Stars
Since the time to merger at detection is a relevant information for multimessenger studies, we also provide the cumulative distributions of the fraction of detected events vs. the time to merger at the moment in which the threshold SNR = 12 is reached. (Figures 14 and 15). At the EMR site, we notice that in the worst possible scenario only $11\%$ of events are detectable within 12 minutes from merger, whereas, in the best noise conditions, this fraction improves to $45\%$. For the 50th percentile case, on the other hand, the fraction is $30\%$. In Sardinia the difference between the three cases is less apparent, since the fraction of events detectable within 12 minutes from merger oscillates between $40\%$ and $47\%$.

### v6

@@ 3.1 Seismic
Figure 3: Seismic ASD displacement for the two candidate sites, TERZ in red and P2 in blue, are reported as a percentile (50th in solid line while 10th and 90th the lower and upper limit of the colored band). Peterson limits (NLNM in black dashed line and NHNM in black dotted line) and the self-noise of both sensors (black solid line) are also shown.
@@
Figure 3 shows the comparison of the percentiles of the horizontal component for P2 (in blue) and TERZ (in red) with respect to the Peterson New Low and High Noise Model (NLNM in black dashed line and NHNM in black dotted line respectively) [89]. Above 6 Hz, the P2 data lie on the self-noise of the sensor, making apparent the quietness of the site. We will use these data in Section 4.1 as input parameters for the determination of the ET sensitivity curve.
3.2 Astrophysical Sources
@@
The parameters for IMBH and BNS mergers are taken from the catalog [90] used to produce the results of [8], which considers a set of sources spanning one year between the mock dates 01-01-2030 and 12-31-2030. For what concerns BBH, the catalog results from the mixture of BBH from isolated binary evolution, dynamical formation in young, globular and nuclear star clusters. Masses, spins, redshifts and luminosity distances have been obtained with the open-source code FASTCLUSTER [91]. For this work, we select BBH with at least one source frame component mass above 100 $M_{\bigodot}$, for a total of $\simeq 10^{3}$ events. Figure 4(a) (blue histogram) shows the distribution of the time to the merger, calculated numerically from the waveforms, at $10\rm\,Hz$ for these events. We note that once left the frequency band considered in this work, most events have less than $0.5\rm\,s$ before merging. On the other hand Figure 4(b) (blue histogram) shows the time spent in the $[2,10]\rm\,Hz$ range and points out how most of the time IMBH lie in this bandwidth before the merger.
(a) Distribution of the time to merger at $10\rm\,Hz$ for the events considered in this work.
@@
(b) Distribution of the time spent in the $[2,10]\rm\,Hz$ range for the events considered in this work.
Figure 4: Times distributions for the events accounted for in this work.
@@
For BNS mergers, the source frame masses of the two objects are sampled uniformly in the interval $[1.1;2.5]\rm M_{\bigodot}$ so that $m_{1}>m_{2}$. Redshifts and luminosity distances are taken from [92]. Since [90] contains $\mathcal{O}(7\times 10^{5})$ BNS events, to limit the computational cost and time of the analysis we randomly sample, using a logarithmic distribution in redshift to privilege closer - therefore stronger - signals, $10^{3}$ events at $z<1.5$. As mentioned, the BNS signal should reach $SNR\geq 12$ to issue an early warning between 20 hours and 1 hour before the merger. Figure 4(a) (orange histogram) shows the distribution of the time to the merger at $10\rm\,Hz$ for the selected BNS events. We note that, once left the [2,10] Hz frequency band, most events have less than $20\rm\,{min}$ before merger. Figure 4(b) (orange histogram) shows the time spent in the $[2,10]\rm\,Hz$ range. The relevance of this bandwidth is apparent, since several hours are spent here.
@@
4 Results
4.1 Modified ET curves
The ET sensitivity, modified according to what described in Section 2, takes into account the local seismic component of the NN of the two candidate sites. In equation 1, we assume $\rho_{0}=2.7\cdot 10^{3}$ kg/m3 and $L=10$ km. Figure 5 reports the sensitivity curve using the NN from P2 and TERZ compared with the ET design noise budget as reported in [8] for the frequency range between 1 Hz and 100 Hz.
@@ 4.1 Modified ET curves
To quantify the modified ET sensitivity with respect to the designed one, we evaluated the NTR index defined in Sec. 2. While Fig. 5 shows the level of the NN inside the ET sensitivity averaged over two years, the NTR, computed every $120\rm\,s$, gives information on how many times the modified sensitivity ($S_{n}(f)$) is above ($>1$) ore below ($<1$) the designed one ($S_{n,\rm{ET}}(f)$), i.e., information related to the glitchiness of the site. Focusing our attention on the frequency range between $f_{1}=2$ Hz and $f_{2}=10$ Hz, a value of NTR larger than one implies that, in the selected dataset, the contribution of the NN is limiting the ET sensitivity. The higher the NTR, the higher the probability of losing the GW signals. One has to remember that if the bandwidth of the GW signal is larger than $f_{2}$ it is possible to recover the signal since the NN is not significant at higher frequencies.
@@
Perc. P2 TERZ P2 TERZ P2 TERZ P2 TERZ P2 TERZ
@@
4.2 Impact on Signal to Noise Ratio
4.2.1 Intermediate Mass Black Holes
Figure 8 shows the distributions of the ratio between the measured SNR and the SNR calculated for the design sensitivity case ($\frac{SNR}{SNR_{\rm{DESIGN}}}$) for the IMBH case for the Sardinia and EMR candidate sites, respectively. These distributions show any loss/gain of SNR with respect to the design sensitivity case. The distributions of the SNR after which the $\frac{SNR}{SNR_{\rm{DESIGN}}}$ ratio is inferred are reported in the Appendix. The overall results, including all classes of astrophysical objects considered in this work, are also summarized in Table 2. For both Sardinia and EMR cases, we analyze the same 970 IMBH events and we exclude all those events which have $\rm SNR<12$ in the best noise condition already. These events never reach the SNR = 12 threshold at 10 Hz also when the noise worsens and therefore are not relevant for the scope of this work. On the contrary, this cut guarantees that events eventually lost in the design case, but recovered in the best noise conditions, are retained. This can also highlight any possible improvement of the performances with respect to the expectations.
In Sardinia, on average, the 10th and 50th percentile cases show comparable results with an SNR gain of $\simeq 7\%$ and $\simeq 6\%$, respectively (Figures 8(a) and 8(b), blue histogram). On the other hand, in the 90th percentile case the performance is on par with the design case: there is only a marginal SNR loss ($\simeq 2.5\%$ Figure 8(c), blue histogram).
For the EMR site, the situation is different. The 10th percentile case is on par with the design performance, with the SNR showing a gain of $\simeq 2.5\%$ (Figure 8(a), orange histogram). In the 50th percentile, the SNR distribution starts do deviate from the design case, with a SNR loss $\simeq 14\%$ (Figure 8(b), orange histogram). The 90th percentile shows, as expected, the worst performance with an average SNR loss $\simeq 37\%$ (Figure8(c), orange histogram). A summary of the results is reported in Table 2.
@@ 4.2.1 Intermediate Mass Black Holes
Perc. Sardinia EMR Sardinia EMR
10th $+7\%$ $+2.5\%$ $-2\%$ $-1\%$ $0.96$
50th $+5\%$ $-13\%$ $-1.8\%$ $5\%$ $0.83$
90th $-2.5\%$ $-37\%$ $2.5\%$ $24\%$ $0.69$
@@
Table 2: Summary of the SNR performance for the Sardinia and EMR candidate sites. The figures reported in this table are the average of the total IMBH and BNS cases. The separate tables for the two classes of events are reported in the Appendix. The negative figures under the events with SNR $<12$ highlight an improvement in the fraction of recovered events with respect to the design case.
(a) Sardinia SNR/SNRdesign using the 10th percentile ET sensitivity.
(b) Sardinia SNR/SNRdesign using the 50th percentile ET sensitivity.
(c) Sardinia SNR/SNRdesign using the 90th percentile ET sensitivity.
Figure 8: IMBH SNR loss distributions for the Sardinia (blue) and EMR (orange) candidate sites for the 10th, 50th and 90th percentile of the seismic noise. The target SNR is shown as a vertical black dashed line.
(a) Sardinia SNR/SNRdesign using the 10th percentile ET sensitivity.
(b) Sardinia SNR/SNRdesign using the 50th percentile ET sensitivity.
(c) Sardinia SNR/SNRdesign using the 90th percentile ET sensitivity.
Figure 9:
BNS SNR loss distributions for the Sardinia (blue) and EMR (orange) candidate sites for the 10th, 50th and 90th percentile of the seismic noise. The target SNR is shown as a vertical black dashed line.
Figure 10: Cumulative distributions of the time to merger at detection for the Sardinia (blue) and EMR (orange) sites.
4.2.2 Binary Neutron Stars
@@
Figure 4.2.1 shows the results obtained in the case of BNS signals. As for the previous results, we exclude events with SNR $<12$ in the best noise case. For Sardinia, the results are consistent with the IMBH analysis and, in all cases, the ratios between the reconstructed SNR and the design SNR show only marginal differences with respect to what discussed in the previous section (Figure 4.2.1, blue histograms). The EMR site shows again the less promising performances. Again (Figure 8, orange histograms) the BNS case show results which are consistent with the IMBH case, with the 10th percentile case equaling the figure of the benchmark case, with a SNR gain $\simeq 2\%$, and the 50th and 90th percentile showing the worst performances loosing $\simeq 14\%$ and $\simeq 37\%$ of the SNR, respectively. Since the time to merger at detection is a relevant information for multimessenger studies, we also provide the cumulative distributions of the fraction of detected events vs. the time to merger at the moment in which the threshold SNR = 12 is reached (Figure 10). The estimation of the time to merger is done numerically from the simulated waveforms, i.e., we mark the time at which each event reaches the SNR threshold and infer from the waveform how far it is from the merger time. At the EMR site, we notice that in the worst possible scenario only $11\%$ of events are detectable within 12 minutes from merger, whereas, in the best noise conditions, this fraction improves to $45\%$. For the 50th percentile case, on the other hand, the fraction is $30\%$. In Sardinia the difference between the three cases is less apparent, since the fraction of events detectable within 12 minutes from merger oscillates between $40\%$ and $47\%$.
5 Discussion
We investigated the impact of site dependent ambient noise on the low-frequency sensitivity of ET at the two sites candidate to host ET: the Sos Enattos site in Sardinia, Italy, and the Terziet side in the EMR region at the border between The Netherlands and Belgium. According to the currently available site characterization data, we focused on the impact of the seismic noise contribution to NN which is more prominent between $2\rm\,Hz$ and $10\rm\,Hz$. Using the 10th, 50th and 90th percentile of the seismic spectra distributions covering two years at the two candidate sites, we inferred the impact on the ET sensitivity curve. We found that the Sardinia candidate site has the lowest impact on the ET sensitivity if compared to the effect of local ambient noise recorded at EMR, which, on the other hand, appears to have a larger impact. These results were obtained without considering any NN suppression factor in Equation 1. As a consequence, we expect that these results may provide some relevant information about the design of NN suppression systems for ET that will be different for each candidate site. In fact, Sardinia appears to be compliant with the requirements of ET already without considering any NN suppression factor, whereas the larger impact on the design sensitivity of the EMR site will require dedicated seismometer arrays for noise subtraction [68] to reduce its impact on the design sensitivity. These results suggests that Sardinia is also an ideal location for research and development activities in low noise environments. For what concerns achievable NN suppression factors, previous works [70, 93, 94] showed that by optimizing NN cancellation for a specific frequency may lead up to a factor 10 suppression. Nevertheless, at other frequencies the factor would be $\simeq 2$ [93]. Broadband NN cancellation, on the other hand, would not achieved a suppression factor better than 3 [93]. It should also be pointed out that the signals considered in this work span a wide range of frequencies, therefore we should focus on broadband NN cancellation. Finally, NN cancellation comes at the cost of a huge computational and financial effort. For example, considering the achievement of a factor 3 NN suppression factor as the most reasonable assumption for ET, it would require a few tens of seismometers in boreholes per test mass [93]. Since the ET triangle will have 12 test masses, we should expect $\mathcal{O}(100)$ sensors placed in boreholes around the test masses. The main challenges would be determining where to drill the boreholes to achieve effective NN reduction and the processing of 3D data for the NN cancellation algorithms (being on surface, current detectors only require 2D data). The different impact of the Sardinia and EMR sites on the ET sensitivity translates into an influence on the signal SNR at low frequency for those astrophysical sources for which low-frequency sensitivity plays an important role, i.e. IMBH and BNS mergers. Using the source parameters present in astrophysical catalogs obtained from population studies, we generated a set of waveforms and injected them into ET noise simulated according to the sensitivity curves derived from local ambient noise. SNR has then been calculated using the matched filter SNR. Comparing the resulting SNR distributions against the benchmark case obtained from the ET design sensitivity curve, we find that local ambient noise has a lower impact on the signal SNR in Sardinia, which is very close to design. The EMR site shows a higher SNR degradation, with respect to design (Figures 8 and 4.2.1).
6 Conclusions
Given the ET early warning requirements for BNS and the time to merger at $10\rm\,Hz$ of these signals, a significant degradation of the signal SNR may affect the ET early warning capabilities and hinder the possibility of extensive multimessenger observation campaigns for a number of events larger than expected. In this sense, we have also inferred the cumulative distributions of the time to merger at detection and found that this is most likely to happen at EMR than Sardinia (Figure 10). These issues could be solved only with the use of appropriate NN suppression that should be more prominent at the EMR site. In future, we will follow up this study to consider the ET 2L configuration and assess the impact of site dependent noise on a detector network composed of two widely separated detectors that are not affected by the same noise sources at the same times. Perhaps, in this case, the presence of the most sensitive interferometer in Sardinia could make the impact of the worse EMR sensitivity on the detection rates less dramatic. We also plan to quantify the effect of local ambient noise on the source sky localization areas issued by early warnings using currently available detection pipelines. This will further expand the overview of the impact of ambient noise on the detectability of GW sources.
The authors gratefully acknowledge dr. M. Mancarella, prof. T. Bulik, dr. S. Koley and dr. J. van Heijningen for their helpful comments and advices that significantly improved the presentation of this work. Dr. M. Di Giovanni would also like to acknowledge dr. E. Codazzo for providing the necessary information to correctly download and interpret the information contained in the astrophysical catalogs. This work is also partially funded by the PRIN 2020 PE9 project 2020BRP57Z of the Ministero dell’Università e della Ricerca.
Appendix
For completeness, in the Appendix we report the SNR distributions for the Sardinia and EMR sites for both the IMBH and BNS cases (Figures A1-A4). These distributions are those used to calculate the $\frac{SNR}{SNR_{\rm DESIGN}}$ ratio of Figures 8 and 4.2.1. In the various figures, the reference distribution of the design sensitivity case is shown in green. As mentioned in the paper, we exclude from the analysis all the events that never reach SNR = 12 below 10 Hz in the 10th percentile case, since they will be never detected in any other case as well. This also implies that some events recovered in the 10th percentile case (best possible noise conditions) may be missed in the benchmark case, since the 10th percentile sensitivity curve may be, at some frequencies, better than the design one. This is indicated by the green histogram bins appearing below the SNR = 12 threshold. We also include the tables (Tables A1 and A2) which contain the results for the IMBH and BNS cases separately that are summarized in Table 2.
(a) Sardinia Nominal SNR distribution (green) and SNR distribution using the 10th percentile ET sensitivity (blue).
(b) Sardinia nominal SNR distribution (green) and SNR distribution using the 50th percentile ET sensitivity (blue).
(c) Sardinia nominal SNR distribution (green) and SNR distribution using the 90th percentile ET sensitivity (blue).
Figure A1: IMBH SNR distributions for the Sardinia candidate site. In the top row, the SNR = 12 threshold is marked with a black dashed line. In the bottom row, the black dashed line represents the equality between the SNR measured in the design case with the cases of the modified ET curves. The non overlapped green SNR bin under the SNR threshold highlights the fact that there are some events lost in the design case but recovered in the best possible recorded noise conditions.
(a) EMR Nominal SNR distribution (green) and SNR distribution using the 10th percentile ET sensitivity (blue).
(b) EMR Nominal SNR distribution (green) and SNR distribution using the 50th percentile ET sensitivity (blue).
(c) EMR Nominal SNR distribution (green) and SNR distribution using the 90th percentile ET sensitivity (blue).
Figure A2: IMBH SNR distributions for the EMR candidate site. In the top row, the SNR = 12 threshold is marked with a black dashed line. In the bottom row, the black dashed line represents the equality between the SNR measured in the design case with the cases of the modified ET curves. The non overlapped green SNR bin under the SNR threshold highlights the fact that there are some events lost in the design case but recovered in the best possible recorded noise conditions.
(a) Sardinia Nominal SNR distribution (green) and SNR distribution using the 10th percentile ET sensitivity (blue).
(b) Sardinia nominal SNR distribution (green) and SNR distribution using the 50th percentile ET sensitivity (blue).
(c) Sardinia nominal SNR distribution (green) and SNR distribution using the 90th percentile ET sensitivity (blue).
Figure A3:
BNS SNR distributions for the Sardinia candidate site. In the top row, the SNR = 12 threshold is marked with a black dashed line. In the bottom row, the black dashed line represents the equality between the SNR measured in the design case with the cases of the modified ET curves. The non overlapped green SNR bin under the SNR threshold highlights the fact that there are some events lost in the design case but recovered in the best possible recorded noise conditions.
(a) EMR Nominal SNR distribution (green) and SNR distribution using the 10th percentile ET sensitivity (blue).
(b) EMR Nominal SNR distribution (green) and SNR distribution using the 50th percentile ET sensitivity (blue).
(c) EMR Nominal SNR distribution (green) and SNR distribution using the 90th percentile ET sensitivity (blue).
Figure A4: BNS SNR distributions for the EMR candidate site. In the top row, the SNR = 12 threshold is marked with a black dashed line. In the bottom row, the black dashed line represents the equality between the SNR measured in the design case with the cases of the modified ET curves. The non overlapped green SNR bin under the SNR threshold highlights the fact that there are some events lost in the design case but recovered in the best possible recorded noise conditions.